OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 5(x + 2) greater than or equal to 7x + 2(x - 1)? A number line with a closed circle on 3 and shading to the right A number line with a closed circle on 3 and shading to the left 3 greater than or equal to x x less than or greater to 3

OpenStudy (radar):

First solve the inequality. 5(x + 2)>=7x + 2(x-1) 5x + 10 >=7x + 2x -2 -4x >=-12 divide by -4 x<=3 Now can you pick out the ones that do NOT described the iequality.